# Unlocking massively parallel spectral proper orthogonal decompositions in the PySPOD package

### v1

@@ 1 Introduction
Over the last several years, many tools have been proposed to mining information from spatio-temporal data, such as the proper orthogonal decomposition (POD) Berkooz et al. (1993); Maulik and Mengaldo (2021), the dynamic mode decomposition (DMD) Schmid (2010), and the spectral proper orthogonal decomposition (SPOD) Lumley (1967); Towne et al. (2018); Schmidt et al. (2019); Schmidt and Colonius (2020); Lario et al. (2022). Yet, the majority of these tools comes with limited parallel capabilities for handling large datasets.
In this paper, we propose a first parallel (distributed) SPOD algorithm, that allows large-scale modal decomposition analyses. The algorithm has been tested on both geophysical and fluid mechanics data, reaching 199 terabytes (TB) of size. Weak and strong scalability have been thoroughly assessed for the entire algorithm as well as for its main constituting steps, including input/output handling, discrete Fourier transform, and eigenvalue computations. The efficiency of the novel parallel algorithm proposed in this paper allows analysis of data at unprecedented scales, opening the path towards uncovering new physics in large datasets, that existing SPOD packages were unable to tackle Schmidt (2022); Mengaldo and Maulik (2021).
@@ 2.2 Theory
For the sake of reader’s convenience, we report here the theory behind SPOD, closely following Towne et al. (2018).
@@
where $\Omega$ denotes the spatial domain, $\boldsymbol{W}(\boldsymbol{x})$ the spatial weighting, and the asterisk superscript represents the conjugate transpose. By invoking the Karhunen-Loéve (KL) theorem Kosambi (2016); Loeve (2017), we know that there exists a set of mutually orthogonal deterministic functions that form a complete basis in $H$. This can be defined as $\hat{\boldsymbol{q}}(\boldsymbol{x},f)=\sum_{k=1}^{\infty}a_{k}(f)\boldsymbol{\phi}_{k}(\boldsymbol{x},f)$, where $\hat{(\cdot)}$ denotes the Fourier transform in time. The eigenfunctions, $\boldsymbol{\phi}_{k}$, and their associated eigenvalues, $\lambda_{k}$, are solutions to the eigenvalue problem in the frequency domain:
@@
where we used equation (2), under the assumption of wide-sense stationarity of the stochastic process $\boldsymbol{q}(\boldsymbol{x},t)$ and $\tau=t-t^{\prime}$ is the difference between the two times $t$ and $t^{\prime}$. This implies that the covariance depends only on the difference between two times, $t$ and $t^{\prime}$ and, therefore, we can write, $\boldsymbol{\mathcal{C}}(\boldsymbol{x},\boldsymbol{x}^{\prime},t,t^{\prime})\rightarrow\boldsymbol{\mathcal{C}}(\boldsymbol{x},\boldsymbol{x}^{\prime},\tau)$. The last step allows reformulating the problem in the spectral (also referred to as frequency) domain, as per equation (6). This significant result was first presented in  Lumley (1967, 2007) and revisited in Towne et al. (2018). It provides eigenmodes at each frequency that inherit the same properties of the more traditional (non-spectral) POD.
The SPOD formulation just summarized leads to monochromatic SPOD modes that optimally characterize the second-order space-time moments of the continuous-time stochastic process considered Towne et al. (2018).
@@ 2.3 Practical implementation
Each data block $\ell$ in equation (10) contains $N_{f}$ time snapshots, overlaps by $N_{\mathrm{overlap}}$ time snapshots with the adjacent block, and is regarded as equally representative of the whole data by the ergodic assumption. Indeed, it is a realization of the stochastic process described by the discrete data $\mathbf{Q}$ in equation (9). This approach of partitioning the time series into overlapping data blocks is the well-known Welch periodogram method Welch (1967); Schmidt and Colonius (2020). In the following, we will use the terms realization and data block interchangeably.
@@
where $L-1$ is a normalization factor, that is only appropriate if the data is centered about the sample mean rather than the long-time (i.e., true) mean. The construction of the cross-spectral density matrix in equation (13) finally allows us to write the discrete analog of the frequency-domain eigenvalue problem defined in equation (6):
@@
The SPOD modes, $\boldsymbol{\Phi}_{k}$, and associated modal energies (or eigenvalues), $\boldsymbol{\Lambda}_{k}$, can be computed by solving equation (14) for each frequency, $f_{k}$. In practice, to alleviate the computational burden of this step, one usually turns to the method of snapshots Sirovich (1987):
$\hat{\mathbf{Q}}^{*}_{k}\mathbf{W}\hat{\mathbf{Q}}_{k}\boldsymbol{\Psi}_{k}=\boldsymbol{\Psi}_{k}\boldsymbol{\Lambda}_{k},\quad\boldsymbol{\Phi}=\hat{\mathbf{Q}}\boldsymbol{\Psi}\boldsymbol{\Lambda}^{-1/2}_{k}.$ (15)
@@
where, for real data, $\mathbf{Q}\in\mathbb{R}^{M\times N_{t}}$, the total number of frequencies is $N_{f/2}=\lceil\frac{N_{f}}{2}\rceil+1$. This is because the transformed data at negative frequencies correspond to the conjugates of the positive frequencies, and it is therefore redundant. For additional details on the SPOD method, the interested reader can refer to Towne et al. (2018); Schmidt and Colonius (2020).
@@ 2.4 SPOD for data compression
As shown in Nekkanti and Schmidt (2021), and further explored in Lario et al. (2022), it is possible to compute a matrix of expansion coefficients $\mathbf{A}$. This can be constructed using a weighted oblique projection of the data onto the modal basis
@@
where $\tilde{\mathbf{Q}}$ is an approximation of the original data $\mathbf{Q}$, given the truncation imposed on the number of SPOD modes and frequencies. The storage of the expansion coefficients and SPOD modes required to reconstruct the high-dimensional data in equation (18), can potentially lead to significant savings in terms of memory storage. In addition, the ability to truncate number of SPOD modes and number of frequencies can be beneficial to just store the frequencies of interest (e.g., removing high-frequency noise from the data), and capture low-rank behaviour of the process under study Towne et al. (2018); Schmidt et al. (2019).
For instance, if we consider a problem constituted by 30,000 time snapshots, 37,000,000 spatial points, and 1 variables, we have: $N_{t}\times M_{\mathrm{space}}\times M_{\mathrm{vars}}=30,000\times 37,000,000\times 1$. To store this dataset in memory, we require approximately $8.88$ TB in double floating-point precision, and $4.44$ TB in single floating-point precision.
However, if we, for instance, store the first 3 SPOD modes and 100 frequencies of interest, the amount of storage memory required is significantly lowered. To store the SPOD modes, we would need $3\times 100\times 37,000,000$, that leads to $0.088$ TB of memory in double floating-point precision, and $0.044$ TB in single floating-point precision. To store the time coefficients, we would need $3\times 100\times 30,000$, that leads to $0.000072$ TB in double floating-point precision and $0.000036$ TB in single floating-point precision. Hence, if we store 3 SPOD modes and 100 frequencies, the storage memory required is only 1% of the memory required to store the original dataset, and the data compression s achieved by having full control on what modes and frequencies are left out (if any).
@@ 3.1 SPOD algorithm
The parallelization strategy for the SPOD algorithm uses a single-program multiple-data (SPMD) approach. It allows maintaining the structure of the code similar to that of the serial one, only introducing parallel communication and synchronization in a limited number of places. For the SPOD algorithm, this consists of decomposing the spatial dimensions $M_{\mathrm{space}}$ (conveniently flattened) of the data matrix in equation (9). This is a practical and advantageous choice as it allows preserving all operations in time – in particular, the DFT – without needing expensive all-to-all communication patterns. Additionally, the number of spatial points $M_{\mathrm{space}}$ is usually much larger than the number of variables $M_{\mathrm{vars}}$, thus allowing for considerably higher parallelism. The decomposition of the space dimension allows a straightforward parallel implementation that only requires one single MPI collective reduction operation. Most of the MPI operations are implemented in an auxiliary utility module. This allows using MPI routines off-the-shelf, requiring minor modifications to the original (serial) code.
Entering into more detail, and assuming that the data distribution is achieved right after I/O, the parallelization of the algorithm becomes trivial. Indeed, it consists of a simple parallel reduction operation for the inner product derived from the method of snapshots in equation (15). The MPI-based implementation uses the mpi4py package Dalcin et al. (2005); Dalcin and Fang (2021). In figure 1, we depict the parallel SPOD algorithm. In red, we represent the parallel operations required, while in blue are operations that remain unchanged compared to the serial code. The six steps reported in figure 1 are described in the following.
@@
    Step 5) Eigen-decomposition (see equation (15)): this operation does not require any parallel handling as there is no manipulation of the distributed spatial dimension. Since the size of the blocks is typically small, eigen-decompositions are cheap to compute. Therefore, we perform this step redundantly in all MPI ranks. If this task ever becomes a computational bottleneck, it would be straightforward to replace with an optimized implementation exploiting the available parallelism.
@@ 3.2.1 Background and related work
Since the early 1990s, I/O has been acknowledged as a significant contributor to parallel application performance Crandall et al. (1995). However, relatively little attention was given to I/O compared to the compute and communication subsystems during that time Del Rosario et al. (1993). Since then, the performance gap between the I/O throughput and the compute capability of supercomputers has only increased, hence exacerbating data movement challenges Liu et al. (2014). A 2012 report titled “Storage challenges at Los Alamos National Lab” noted that parallel file systems lack large parallel I/O, which would be both high bandwidth and resistant to I/O patterns without additional tuning Bent et al. (2012). Consequently, I/O systems remain relatively slow and complex, making it difficult for non-expert users to utilize them fully.
@@
Middleware libraries were introduced to improve usability and simplify parallel I/O. After all, as Liu et al. (2014) noted, users need not be aware of their data’s low-level layout and organization, which opens possibilities for specialized I/O middleware. MPI I/O enables simple N–1 I/O ($N$ processes, $1$ file) Thakur et al. (1999), and it has been an ongoing effort since 1999. It contains optimizations reducing the number of distinct I/O requests using techniques such as data sieving and collective I/O for noncontiguous accesses Thakur et al. (2002). MPI I/O, however, is still low-level, as the developer needs to calculate the offsets within the file, while MPI aims to access the data efficiently. Higher-level libraries such as HDF5 The HDF Group (2010), Parallel netCDF Li et al. (2003), ADIOS Liu et al. (2014), and ADIOS 2 Godoy et al. (2020) have also been introduced to isolate developers from low-level file system details and alleviate the burden of platform-specific performance optimization. These libraries often build on top of MPI I/O and introduce custom file formats.
@@
Although robust solutions for parallel I/O are available, achieving good performance often requires application-specific and system-specific tuning. Researchers have dedicated significant effort to optimizing applications that utilize the popular Lustre file system. Tools such as Lustre IO Profiler Xu et al. (2016) were introduced to monitor and understand the I/O activities on the file system. The authors of this tool highlight the importance of incorporating system-specific details, such as the number of Object Storage Targets (OSTs), into the MPI distribution used in order to improve the achieved I/O bandwidth. The authors of Y-Lib Dickens and Logan (2010) also point out the low performance of MPI-IO on Lustre file systems and propose a solution that minimizes contention for file system resources by controlling the number of OSTs with which each aggregator process communicates. This solution is shown to outperform MPI-IO in many cases. Similarly, the authors of Balle and Johnsen (2016) demonstrate that Parallel netCDF with MPI-IO does improve the performance over netCDF; however, even with supercomputer-specific optimizations, the performance still disappoints once a large number of MPI ranks is used. The authors utilize asynchronous I/O (quilt) servers alongside Parallel netCDF to further improve the effective bandwidth. Other solutions explore tuning the middleware or file system parameters, such as those presented in Yu et al. (2008); Howison et al. (2010); Behzad et al. (2013, 2014). These approaches often use sophisticated techniques, such as genetic algorithms, to identify and tune parameters of the I/O stack (HDF5, MPI-IO, and Lustre/GPFS parameters) Behzad et al. (2013, 2014). The wealth of work on the topic highlights the difficulty of achieving good performance on current parallel file systems.
@@
Despite the availability of powerful solutions and tools to tune the performance, many developers still choose not to use them, even in leadership scale systems Xu et al. (2016). An analysis conducted in 2015 using Darshan Luu et al. (2015) found that, depending on the supercomputing facility, between 50 to 95% of monitored jobs used POSIX I/O exclusively. However, as Luu et al. (2015) notes, relying on POSIX I/O does have to result in poor application performance. In fact, in some cases, using a naive N–N ($N$ processes, $N$ files) approach, where each process reads from or writes to a unique file, can achieve better performance and scalability than N–1, as it does not incur overhead to maintain data consistency. This has been observed in PanFS, GPFS, and Lustre file systems. However, the N–N approach presents challenges related to usability, such as when restarting with a different number of processes, and high metadata costs at scale Bent et al. (2012); Yu et al. (2012).

### v2

@@ 1 Introduction
Over the last several years, many tools have been proposed to mining information from spatio-temporal data, such as the proper orthogonal decomposition (POD) [1, 2], the dynamic mode decomposition (DMD) [3], and the spectral proper orthogonal decomposition (SPOD) [4, 5, 6, 7, 8]. Yet, the majority of these tools comes with limited parallel capabilities for handling large datasets.
In this paper, we propose a first parallel (distributed) SPOD algorithm, that allows large-scale modal decomposition analyses. The algorithm has been tested on both geophysical and fluid mechanics data, reaching 199 terabytes (TB) of size. Weak and strong scalability have been thoroughly assessed for the entire algorithm as well as for its main constituting steps, including input/output handling, discrete Fourier transform, and eigenvalue computations. The efficiency of the novel parallel algorithm proposed in this paper allows analysis of data at unprecedented scales, opening the path towards uncovering new physics in large datasets, that existing SPOD packages were unable to tackle [9, 10].
@@ 2.2 Theory
For the sake of reader’s convenience, we report here the theory behind SPOD, closely following [5].
@@
where $\Omega$ denotes the spatial domain, $\boldsymbol{W}(\boldsymbol{x})$ the spatial weighting, and the asterisk superscript represents the conjugate transpose. By invoking the Karhunen-Loéve (KL) theorem [11, 12], we know that there exists a set of mutually orthogonal deterministic functions that form a complete basis in $H$. This can be defined as $\hat{\boldsymbol{q}}(\boldsymbol{x},f)=\sum_{k=1}^{\infty}a_{k}(f)\boldsymbol{\phi}_{k}(\boldsymbol{x},f)$, where $\hat{(\cdot)}$ denotes the Fourier transform in time. The eigenfunctions, $\boldsymbol{\phi}_{k}$, and their associated eigenvalues, $\lambda_{k}$, are solutions to the eigenvalue problem in the frequency domain:
@@
where we used equation (2), under the assumption of wide-sense stationarity of the stochastic process $\boldsymbol{q}(\boldsymbol{x},t)$ and $\tau=t-t^{\prime}$ is the difference between the two times $t$ and $t^{\prime}$. This implies that the covariance depends only on the difference between two times, $t$ and $t^{\prime}$ and, therefore, we can write, $\boldsymbol{\mathcal{C}}(\boldsymbol{x},\boldsymbol{x}^{\prime},t,t^{\prime})\rightarrow\boldsymbol{\mathcal{C}}(\boldsymbol{x},\boldsymbol{x}^{\prime},\tau)$. The last step allows reformulating the problem in the spectral (also referred to as frequency) domain, as per equation (6). This significant result was first presented in  [4, 13] and revisited in [5]. It provides eigenmodes at each frequency that inherit the same properties of the more traditional (non-spectral) POD.
The SPOD formulation just summarized leads to monochromatic SPOD modes that optimally characterize the second-order space-time moments of the continuous-time stochastic process considered [5].
@@ 2.3 Practical implementation
Each data block $\ell$ in equation (10) contains $N_{f}$ time snapshots, overlaps by $N_{\mathrm{overlap}}$ time snapshots with the adjacent block, and is regarded as equally representative of the whole data by the ergodic assumption. Indeed, it is a realization of the stochastic process described by the discrete data $\mathbf{Q}$ in equation (9). This approach of partitioning the time series into overlapping data blocks is the well-known Welch periodogram method [14, 7]. A 50% overlap, i.e., $N_{\mathrm{overlap}}=N_{f}/2$, is accepted best practice. In the following, we will use the terms realization and data block interchangeably.
@@
where $L-1$ is a normalization factor known as Bessel’s correction, that is only appropriate if the data is centered about the sample mean rather than the long-time (i.e., true) mean. The construction of the cross-spectral density matrix in equation (13) finally allows us to write the discrete analog of the frequency-domain eigenvalue problem defined in equation (6):
@@
The SPOD modes, $\boldsymbol{\Phi}_{k}$, and associated modal energies (or eigenvalues), $\boldsymbol{\Lambda}_{k}$, can be computed by solving equation (14) for each frequency, $f_{k}$. In practice, to alleviate the computational burden of this step, one usually turns to the method of snapshots [15]:
$\hat{\mathbf{Q}}^{*}_{k}\mathbf{W}\hat{\mathbf{Q}}_{k}\boldsymbol{\Psi}_{k}=\boldsymbol{\Psi}_{k}\boldsymbol{\Lambda}_{k},\quad{\color[rgb]{0,0,0}\boldsymbol{\Phi}_{k}=\hat{\mathbf{Q}}_{k}\boldsymbol{\Psi}_{k}\boldsymbol{\Lambda}^{-1/2}_{k}}.$ (15)
@@
where, for real data, $\mathbf{Q}\in\mathbb{R}^{M\times N_{t}}$, the total number of frequencies is $N_{f/2}=\lceil\frac{N_{f}}{2}\rceil+1$. This is because the transformed data at negative frequencies correspond to the conjugates of the positive frequencies, and it is therefore redundant. For additional details on the SPOD method, the interested reader can refer to [5, 7].
@@ 2.4 SPOD for data compression
As shown in [16], and further explored in [8], it is possible to compute a matrix of expansion coefficients $\mathbf{A}$. This can be constructed using a weighted oblique projection of the data onto the modal basis
@@
where $\tilde{\mathbf{Q}}$ is an approximation of the original data $\mathbf{Q}$, given the truncation imposed on the number of SPOD modes and frequencies. The storage of the expansion coefficients and SPOD modes required to reconstruct the high-dimensional data in equation (18), can potentially lead to significant savings in terms of memory storage. In addition, the ability to truncate number of SPOD modes and number of frequencies can be beneficial to just store the frequencies of interest (e.g., removing high-frequency noise from the data), and capture low-rank behaviour of the process under study [5, 6].
For instance, if we consider a problem constituted by 30,000 time snapshots, 37,000,000 spatial points, and 1 variable, we have: $N_{t}\times M_{\mathrm{space}}\times M_{\mathrm{vars}}=30,000\times 37,000,000\times 1$. To store this dataset in memory, we require approximately $8.88$ TB in double floating-point precision, and $4.44$ TB in single floating-point precision.
However, if we, for instance, store the first 3 SPOD modes and 100 frequencies of interest, the amount of storage memory required is significantly lowered. To store the SPOD modes, we would need $3\times 100\times 37,000,000\times 2$, that leads to $0.18$ TB of memory in double floating-point precision, and $0.089$ TB in single floating-point precision. To store the time coefficients, we would need $3\times 100\times 30,000\times 2$, that leads to $0.00014$ TB in double floating-point precision and $0.000072$ TB in single floating-point precision. Hence, if we store 3 SPOD modes and 100 frequencies, the storage memory required is only 2% of the memory required to store the original dataset, and the data compression is achieved by having full control on what modes and frequencies are left out (if any).
@@ 3.1 SPOD algorithm
The parallelization strategy for the SPOD algorithm uses a single-program multiple-data (SPMD) approach (the interested reader can also refer to other parallelized modal decompositions, such as [17], that is specialized to boundary-layer flows). It allows maintaining the structure of the code similar to that of the serial one, only introducing parallel communication and synchronization in a limited number of places. For the SPOD algorithm, this consists of decomposing the spatial dimensions $M_{\mathrm{space}}$ (conveniently flattened) of the data matrix in equation (9). This is a practical and advantageous choice as it allows preserving all operations in time – in particular, the DFT – without needing expensive all-to-all communication patterns. Additionally, the number of spatial points $M_{\mathrm{space}}$ is usually much larger than the number of variables $M_{\mathrm{vars}}$, thus allowing for considerably higher parallelism. The decomposition of the space dimension allows a straightforward parallel implementation that only requires one single MPI collective reduction operation. Most of the MPI operations are implemented in an auxiliary utility module. This allows using MPI routines off-the-shelf, requiring minor modifications to the original (serial) code.
Entering into more detail, and assuming that the data distribution is achieved right after I/O, the parallelization of the algorithm becomes trivial. Indeed, it consists of a simple parallel reduction operation for the inner product derived from the method of snapshots in equation (15). The MPI-based implementation uses the mpi4py package [18, 19]. In figure 1, we depict the parallel SPOD algorithm. In red, we represent the parallel operations required, while in blue are operations that remain unchanged compared to the serial code. The six steps reported in figure 1 are described in the following.
@@
    Step 5) Eigen-decomposition (see equation (15)): this operation does not require any parallel handling as there is no manipulation of the distributed spatial dimension. Since the size of the blocks is typically small, eigen-decompositions are cheap to compute. Therefore, we perform this step redundantly in all MPI ranks. If this task ever becomes a computational bottleneck, it would be straightforward to replace with a distributed-memory implementation exploiting the already available CPU resources, or alternatively, GPU off-loading for acceleration.
@@ 3.2.1 Background and related work
Since the early 1990s, I/O has been acknowledged as a significant contributor to parallel application performance [20]. However, relatively little attention was given to I/O compared to the compute and communication subsystems during that time [21]. Since then, the performance gap between the I/O throughput and the compute capability of supercomputers has only increased, hence exacerbating data movement challenges [22]. A 2012 report titled “Storage challenges at Los Alamos National Lab” noted that parallel file systems lack large parallel I/O, which would be both high bandwidth and resistant to I/O patterns without additional tuning [23]. Consequently, I/O systems remain relatively slow and complex, making it difficult for non-expert users to utilize them fully.
@@
Middleware libraries were introduced to improve usability and simplify parallel I/O. After all, as [22] noted, users need not be aware of their data’s low-level layout and organization, which opens possibilities for specialized I/O middleware. MPI-IO enables simple N–1 I/O ($N$ processes, $1$ file) [24], and it has been an ongoing effort since 1999. It contains optimizations reducing the number of distinct I/O requests using techniques such as data sieving and collective I/O for noncontiguous accesses [25]. MPI-IO, however, is still low-level, as the developer needs to calculate the offsets within the file, while MPI aims to access the data efficiently. Higher-level libraries such as HDF5 [26], Parallel netCDF [27], ADIOS [22], and ADIOS 2 [28] have also been introduced to isolate developers from low-level file system details and alleviate the burden of platform-specific performance optimization. These libraries often build on top of MPI-IO and introduce custom file formats.
@@
Although robust solutions for parallel I/O are available, achieving good performance often requires application-specific and system-specific tuning. Researchers have dedicated significant effort to optimizing applications that utilize the popular Lustre file system. Tools such as Lustre IO Profiler [29] were introduced to monitor and understand the I/O activities on the file system. The authors of this tool highlight the importance of incorporating system-specific details, such as the number of Object Storage Targets (OSTs), into the MPI distribution used in order to improve the achieved I/O bandwidth. The authors of Y-Lib [30] also point out the low performance of MPI-IO on Lustre file systems and propose a solution that minimizes contention for file system resources by controlling the number of OSTs with which each aggregator process communicates. This solution is shown to outperform MPI-IO in many cases. Similarly, the authors of [31] demonstrate that Parallel netCDF with MPI-IO does improve the performance over netCDF; however, even with supercomputer-specific optimizations, the performance still disappoints once a large number of MPI ranks is used. The authors utilize asynchronous I/O (quilt) servers alongside Parallel netCDF to further improve the effective bandwidth. Other solutions explore tuning the middleware or file system parameters, such as those presented in [32, 33, 34, 35]. These approaches often use sophisticated techniques, such as genetic algorithms, to identify and tune parameters of the I/O stack (HDF5, MPI-IO, and Lustre/GPFS parameters) [34, 35]. The wealth of work on the topic highlights the difficulty of achieving good performance on current parallel file systems.
@@
Despite the availability of powerful solutions and tools to tune the performance, many developers still choose not to use them, even in leadership scale systems [29]. An analysis conducted in 2015 using Darshan [36] found that, depending on the supercomputing facility, between 50 to 95% of monitored jobs used POSIX I/O exclusively. However, as [36] notes, relying on POSIX I/O does have to result in poor application performance. In fact, in some cases, using a naive N–N ($N$ processes, $N$ files) approach, where each process reads from or writes to a unique file, can achieve better performance and scalability than N–1, as it does not incur overhead to maintain data consistency. This has been observed in PanFS, GPFS, and Lustre file systems. However, the N–N approach presents challenges related to usability, such as when restarting with a different number of processes, and high metadata costs at scale [23, 37].